OpenStudy (anonymous):

Prove directly from the definition that: lim n^2/(n^(2)+1) = 1

OpenStudy (campbell_st):

you need to prove \[\frac{\lim_{n \rightarrow ?} n^2}{\lim_{n \rightarrow ?} n^2 + 1}\]

OpenStudy (anonymous):

I am not good with proofs, so I am having a really hard time with this. Could you guide me?

OpenStudy (campbell_st):

so you need to look at \[\frac{(n^2 +1) -1}{n^2 + 1} \] which is tha same as \[\frac{n^2 +1}{n^2 +1} - \frac{1}{n^2 + 1} \] which simplifies to \[1 - \frac{1}{n^2 + 1}\] which will allow you to find the limit...

OpenStudy (anonymous):

Ok...so where do I go from here?

OpenStudy (anonymous):

I see how you got that. We did a problem similar to this in class, so that part clicked,but we never finished it. So this is where I am stuck even more lol

hero (hero):

What does the limit go to?

OpenStudy (anonymous):

it doesn't specify

OpenStudy (anonymous):

i think infinity though

OpenStudy (anonymous):

it must be infinity, otherwise this is never 1

OpenStudy (anonymous):

thought so

hero (hero):

Okay, so use the definition of limit to solve

OpenStudy (anonymous):

you can also divide top and bottom by \(n^2\) to get the same result, although @campbell_st method is more elegant

OpenStudy (anonymous):

ok, so i am suppose to show my work for the definition of limit to solve but idk how to do that, because I just know how to look at it and just KNOW, i dont know how i am suppose to prove

OpenStudy (helder_edwin):

this is the limit of a squence. right? \[ \large \lim_{n\to\infty}\frac{n^2}{n^2+1}=1 \]

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

let's remember the definition. left \(a_n=n^2/(n^2+1)\) then \[ \large \lim a_n=1 \] if and only if \[ \large (\forall\varepsilon>0)(\exists N)(n\geq N\to|a_n-1|<\varepsilon) \] ok?

OpenStudy (helder_edwin):

let \(\varepsilon>0\) then \[ \large |a_n-1|=\left|\frac{n^2}{n^2+1}-1\right|= \left|\frac{-1}{n^2+1}\right|=\frac{1}{n^2+1}< \varepsilon \] then \[ \large \frac{1}{\varepsilon}<n^2+1 \] \[ \large \frac{1}{\varepsilon}-1<n^2 \] \[ \large n>\sqrt{\frac{1}{\varepsilon}-1} \]

OpenStudy (helder_edwin):

so let \[ \large N=\sqrt{\frac{1}{\varepsilon}-1} \]